Start learning →

P(at least one head in 2 fair coin tosses):

A1/4
B1/2
C3/4
D1
Answer & Solution
Correct answer: C. 3/4
P(no heads) = (1/2) × (1/2) = 1/4. P(at least one head) = 1 - 1/4 = 3/4.
Practise this question with Quill — AP Intermediate 2nd Year →

Keep practising Statistics and Probability

Browse all Statistics and Probability questions for AP Intermediate 2nd Year, or explore more Mathematics topics.

Start a Statistics and Probability lesson →